If we have a random variable T distributed according to exponential distribution with mean 0.8 hours. What is the probability that T will be greater than 48 minutes?
Answer #1

Let, T folloes exponentail distribution with mean = 0.8

We know that,

for exponentail distribtiion P(X<x) = 1- e-x/ α

We need tocompute,

P(T > 48) = 1-P(T<48) = 1-(1-e-48/ 0.8)

P(T > 48) = e-48/ 0.8

P(T > 48) = 0